# Planar percolation and the loop $\mathrm{O}(n)$ model

### v2

@@ 1.3. Loop $\mathrm{O}(n)$ model
Among particular cases of the loop $\mathrm{O}(n)$ model are the Ising model ($n=1$), $1/2$ Bernoulli site percolation ($n=x=1$), dimer model ($n=1,x=\infty$), integer-valued 1-Lipschitz function ($n=2$); see Fig. 2. There are heuristic connections to the spin $\mathrm{O}(n)$ model when $n$ is integer; see [PS19] for a survey.
The above definition can be extended to the infinite volume in a standard way using the formalism Dobrushin, Landford and Ruelle (DLR); see [GL25a] for more details. We endow the set of loop configurations $\mathfrak{S}_{{\sf Loop}}(\mathbb{H})$ with the $\sigma$-algebra generated by cylinder events. A measure $\mu$ on this space is a Gibbs measure if, for any domain $\Omega$ and for $\mu$-almost every $\omega^{\prime}$}, the measure $\mu$ conditional on $\{\omega\in\mathfrak{S}_{{\sf Loop}}(\mathbb{H};\Omega;{\omega^{\prime}})\}$ equals ${\sf Loop}_{\Omega,n,x}^{\omega^{\prime}}$.
A Gibbs measure $\mu$ is called translation invariant if
@@
  • At $x=x_{c}(n)$ when $n\in[1,2]$ [DGPS21],
  • An area $n\in[1,1+\varepsilon],x\in[1-\varepsilon,1/\sqrt{n}]$ around $n=x=1$ [CGHP25],
  • $n=2,x\in[1/\sqrt{2},1]$ [GM21, GL25a].
In addition, this is standard for the loop $\mathrm{O}(1)$ model with $x\in[1/\sqrt{3},1]$ which corresponds to the supercritical Ising model on the triangular lattice (eg. follows from [Tas16]). We also mention that it has been shown that any translation-invariant Gibbs measure for the loop $\mathrm{O}(1)$ model at $x=\infty$ (fully-packed case, complement dimers on the hexagonal lattice) either exhibits infinitely many loops around every face or a unique bi-infinite path [GR24]. The macroscopic behavior is expected to hold for all $n\in[0,2]$ when $x\geq x_{c}(n)$, where
@@
This value was derived first at $n=2$ [DMNS81, Fan72] based on a relation with the Ashkin–Teller model and then for all $n\in[0,2]$ [Nie82] using the Coulomb gas formalism.
The macroscopic behavior is a weaker version of scale invariance and is in agreement with the conjectured convergence to the conformal loop ensemble (CLE) with an appropriate $\kappa=\kappa(n,x)$; see [KN04, Section 5.6]. The convergence has been proved only for Bernoulli site percolation (n=x=1n=x=1[Smi01b, Smi01a, CN06, KS21] and for the critical Ising model ($n=1,\,x=1/{\sqrt{3}}$[Smi10, CS11, CDH+14].
When $n=2$, the loops are level lines of an integer-valued 1-Lipschitz function on the faces of $\mathbb{H}$. Theorem 2 provides a new proof of its delocalization for all $x\in[x_{c}(2),1]$. The localization is expected for all $x<x_{c}(2)$; this has been proven for $x<x_{c}(1)+\varepsilon$ (note that $x_{c}(1)=1/\sqrt{3}$ and $x_{c}(2)=1/\sqrt{2}$).
Besides the use of Theorem 1, the proof of Theorem 2 combines elements of the geometric representation in [GL25a], the ‘flipping’ transformation in [GM21], and the defect percolation construction of [CGHP25]. This approach rules out bi-infinite paths even outside the positive association regime. Inside the positive association regime, the dichotomy theorem of [DGPS21] allows us to upgrade the infinite loop state to a RSW-type result. We mention that [KST23] proves RSW estimates for general positively correlated percolation processes on lattice with $\mathbb{Z}^{2}$ symmetries. Our approach does not rely on this work, and we do not believe it can be used to imply the conclusions of Theorem 2.
@@ 1.4. Sketch of the proof of Theorem 1
The main step of the proof is splitting $\partial\Omega_{n}$ into two arcs, each connected to infinity in $\Omega_{n}^{c}$ with probability almost one (a similar idea was also used in [Tas16]). Denote the vertices on $\partial\Omega_{n}$ by $v_{1},\dots,v_{L}$, taken in clockwise order. For each $i,j\in\{1,\dots,L\}$, let $p_{i,j}$ denote the probability that $\{v_{i},v_{i+1},\dots,v_{j}\}$ does not connect to infinity in $\Omega_{n}^{c}$. Clearly, $p_{1,i}$ is decreasing and $p_{1,L}<\varepsilon$. Define
@@ Acknowledgements
We thank Diederik van Engelenburg, Russell Lyons, and Gabor Pete for insightful discussions on unimodular graphs. The project started from an open problem session at a conference at the ETH Zurich in 2024, and we thank Hugo Duminil-Copin and Vincent Tassion for organizing this wonderful event. Parts of the project were completed during the visits of AG and MH to Northeastern University and the University of Innsbruck, and we would like to thank these institutions for their hospitality.
@@ 2. Planar percolation models: proof of Theorem 1
Let $G=(V,E)$ be a planar, connected, locally finite graph. We will also assume $G$ is simple, as multiple edges and self-loops do not affect site percolation. Given an embedding $\phi:V\to\mathbb{S}^{2}$, we call $a\in\mathbb{S}^{2}$ an accumulation point if its arbitrarily small neighborhood intersects infinitely many edges of $G$. Denote the set of accumulation points by $A$. We call $\phi$ well-separated if $\phi(G)$ is disjoint from $A$. In Proposition 2.2 below, we show that $G$ always has a well-separated embedding.
@@
For $m>n$, define $C_{n,m}(\sigma)$ (resp. $C_{n,m}(1-\tau)$) as the number of connected components of $\Omega_{m}\cap\{\sigma=1\}$ (resp. $\Omega_{m}\cap\{\tau=0\}$) which contain a crossing from $\mathcal{S}_{n}$ to $\mathcal{S}_{m}$ in the annulus domain $\Omega_{m}\smallsetminus\Omega_{n}$. It is enough to show that $C_{n,m}(\sigma)+C_{n,m}(1-\tau)\geq k+1$ for each $m>n$. We prove this by induction on $k$.
@@
Both Corollaries 1.1 and 1.2 follow from Theorem 1 if one can rule out the existence of infinitely many infinite connected components. A beautiful argument of Burton and Keane [BK89] treats a large class of bond percolations on transitive amenable graphs. Here we adapt it to the setting of site percolations on unimodular invariantly amenable graphs. Note that a site percolation might not have any trifurcation point due to combinatorial constraints (e.g. the Kagome lattice). In order to resolve this issue, we sample an auxiliary uniform spanning forest $\tau$ on top of our site percolation $\sigma$ and show that $\tau$ does have trifurcation points (this is reminiscent to [LP17, Lemma $7.7$]). We then extend the argument to unimodular graphs using (MTP) in a fairly standard way. We provide all details of the proof of Corollary 1.1 and then explain how to modify it to the more general setting of Corollary 1.2.
@@
For $n\in\mathbb{N}$, consider the subgraph of $G$ induced by $\{\sigma=1\}\cap\Lambda_{n}(\rho)$ and denote its connected component containing $\rho$ by $\mathcal{C}_{n,\rho}$. Define $\mathsf{UST}_{n,G,\sigma}$ as the uniform spanning tree measure on $\mathcal{C}_{n,\rho}$. It is standard that the UST satisfies a negative correlation inequality, whence the limit of $\mathsf{UST}_{n,G,\sigma}$ as $n$ tends to infinity exists; see [BLPS01]. The limit is called free uniform spanning forest, we denote it by $\mathsf{FUSF}_{G,\sigma}$. Below we do not rely on [AHNR18, Theorem $5.13$] establishing that the FUSF is connected a.s. since that work assumes that $\rho$ has a finite expected degree.
@@
Take any finitary invariant bond percolation $\omega$ and let $T(G,\rho,\omega)$ be the number of trifurcation points of $\tau$ in $K_{\omega}(\rho)$. By the standard 3-partition argument of Burton and Keane (see [Gri99, Lemma 8.5]), $T(G,\rho,\omega)$ is bounded above by $|\partial K_{\omega}(\rho)|$. Now set
@@ 4. Loop $\mathrm{O}(n)$ model: proof of Theorem 2
The next graphical representation is inspired by the Edwards–Sokal coupling [ES88] and the work of Chayes and Machta [CM97]. In a particular case of $n=2$, this representation was introduced by the second author and Spinka; it is described in [GL25a].
@@
Let $n\in[1,2]$ and $x\in[1/\sqrt{2},1]$. Consider any translation invariant Gibbs measure $\mathbb{P}$ for the loop $\mathrm{O}(n)$ model with edge-weight $x$. Sample a loop configuration $\omega$ from $\mathbb{P}$ and then $\xi$ from $\omega$. Then, $\xi$ exhibits no infinite connected component, almost surely.

### v3

@@ 1.3. Loop $\mathrm{O}(n)$ model
Among particular cases of the loop $\mathrm{O}(n)$ model are the Ising model ($n=1$), $1/2$ Bernoulli site percolation ($n=x=1$), dimer model ($n=1,x=\infty$), integer-valued 1-Lipschitz function ($n=2$); see Fig. 2. There are heuristic connections to the spin $\mathrm{O}(n)$ model when $n$ is integer; see [50] for a survey.
The above definition can be extended to the infinite volume in a standard way using the Dobrushin, Landford and Ruelle (DLR) formalism; see [27] for more details. We endow the set of loop configurations $\mathfrak{S}_{{\sf Loop}}(\mathbb{H})$ with the $\sigma$-algebra generated by cylinder events. A measure $\mu$ on this space is a Gibbs measure if, for any domain $\Omega$ and for $\mu$-almost every $\omega^{\prime}$}, the measure $\mu$ conditional on $\{\omega\in\mathfrak{S}_{{\sf Loop}}(\mathbb{H};\Omega;{\omega^{\prime}})\}$ equals ${\sf Loop}_{\Omega,n,x}^{\omega^{\prime}}$.
A Gibbs measure $\mu$ is called translation-invariant if
@@
  • At $x=x_{c}(n)$ when $n\in[1,2]$ [20],
  • An area $n\in[1,1+\varepsilon],x\in[1-\varepsilon,1/\sqrt{n}]$ around $n=x=1$ [18],
  • $n=2,x\in[1/\sqrt{2},1]$ [28, 27].
In addition, this is standard for the loop $\mathrm{O}(1)$ model with $x\in[1/\sqrt{3},1]$ which corresponds to the Ising model on the triangular lattice at $T>T_{c}$ (eg. follows from [57]). We also mention that it has been shown that any translation-invariant Gibbs measure for the loop $\mathrm{O}(1)$ model at $x=\infty$ (fully-packed case, whose complement is uniform dimers on the hexagonal lattice) either exhibits infinitely many loops around every face or a unique bi-infinite path [29]. The macroscopic behavior is expected to hold for all $n\in[0,2]$ when $x\geq x_{c}(n)$, where
@@
This value was derived first at $n=2$ [19, 25] based on a relation with the Ashkin–Teller model and then for all $n\in[0,2]$ [49] using the Coulomb gas formalism.
The macroscopic behavior is a weaker version of scale invariance and is in agreement with the conjectured convergence to the conformal loop ensemble (CLE) with an appropriate $\kappa=\kappa(n,x)$; see [37, Section 5.6]. The convergence has been proved only for Bernoulli site percolation (n=x=1n=x=1[53, 54, 13, 40] and for the critical Ising model ($n=1,\,x=1/{\sqrt{3}}$[55, 17, 16].
When $n=2$, the loops are level lines of an integer-valued 1-Lipschitz function on the faces of $\mathbb{H}$. Theorem 2 provides a new proof of its delocalization for all $x\in[x_{c}(2),1]$. Localization is expected for all $x<x_{c}(2)$; this has been proven for $x<x_{c}(1)+\varepsilon$ (note that $x_{c}(1)=1/\sqrt{3}$ and $x_{c}(2)=1/\sqrt{2}$).
Besides the use of Theorem 1, the proof of Theorem 2 combines elements of the geometric representation in [27], the ‘flipping’ transformation in [28], and the defect percolation construction of [18]. This approach rules out bi-infinite paths even outside the positive association regime. Inside the positive association regime, the dichotomy theorem of [20] allows us to upgrade the infinite loop state to a RSW-type result. We mention that [41] proves RSW estimates for general positively correlated percolation processes on lattice with $\mathbb{Z}^{2}$ symmetries. Our approach does not rely on this work, and we do not believe it can be used to imply the conclusions of Theorem 2.
@@ 1.4. Sketch of the proof of Theorem 1
The main step of the proof is splitting $\partial\Omega_{n}$ into several arcs, each connected to infinity in $\Omega_{n}^{c}$ with probability almost one. Similar approaches appeared in the context of Voronoi percolation on $\mathbb{R}^{2}$ [57], and in level-set percolations of one-Lipschitz functions on cubic, vertex-transitive graphs [38].
Denote the vertices on $\partial\Omega_{n}$ by $v_{1},\dots,v_{L}$, taken in clockwise order. For each $i,j\in\{1,\dots,L\}$, let $p_{i,j}$ denote the probability that $\{v_{i},v_{i+1},\dots,v_{j}\}$ does not connect to infinity in $\Omega_{n}^{c}$. Clearly, $p_{1,i}$ is decreasing and $p_{1,L}<\varepsilon$. Define
@@ Acknowledgements
The project was inspired by a discussion between the second author and Ron Peled about no-coexistence in non translation-invariant percolation processes on $\mathbb{Z}^{2}$. We thank him for sharing his insights. We also thank Diederik van Engelenburg, Russell Lyons, and Gabor Pete for fruitful discussions on unimodular graphs. This collaboration started during an open problem session at a conference at the ETH Zurich in 2024, and we thank Hugo Duminil-Copin and Vincent Tassion for organizing this wonderful event. Parts of the project were completed during the visits of AG and MH to Northeastern University and the University of Innsbruck, and we would like to thank these institutions for their hospitality.
@@ 2. Planar percolation models: proof of Theorem 1
Let $G=(V,E)$ be a planar, connected, locally finite graph. We will also assume $G$ is simple, as multiple edges and self-loops do not affect site percolation. Given an embedding $\phi:V\to\mathbb{S}^{2}$, we call $a\in\mathbb{S}^{2}$ an accumulation point if its arbitrarily small neighborhoods intersect infinitely many edges of $G$. Denote the set of accumulation points by $A$. We call $\phi$ well-separated if $\phi(G)$ is disjoint from $A$. In Proposition 2.2 below, we show that $G$ always has a well-separated embedding.
@@
For $m>n$, define $C_{n,m}(\sigma)$ (resp. $C_{n,m}(1-\tau)$) as the number of connected components of $\Omega_{m}\cap\{\sigma=1\}$ (resp. $\Omega_{m}\cap\{\tau=0\}$) which contain a crossing from $\mathcal{S}_{n}$ to $\mathcal{S}_{m}$ in the annular domain $\Omega_{m}\smallsetminus\Omega_{n}$. It is enough to show that $C_{n,m}(\sigma)+C_{n,m}(1-\tau)\geq k+1$ for each $m>n$. We prove this by induction on $k$.
@@
Both Corollaries 1.1 and 1.2 follow from Theorem 1 if one can rule out the existence of infinitely many infinite connected components. A beautiful argument of Burton and Keane [12] treats a large class of bond percolations on transitive amenable graphs. Here we adapt it to the setting of site percolations on unimodular invariantly amenable graphs. Note that a site percolation might not have any trifurcation point due to combinatorial constraints (e.g. the Kagome lattice). In order to resolve this issue, we sample an auxiliary uniform spanning forest $\tau$ on top of our site percolation $\sigma$ and show that $\tau$ does have trifurcation points (this is reminiscent to [46, Lemma $7.7$]). We then extend the argument to unimodular graphs using (MTP) in a fairly standard way. We provide all details of the proof of Corollary 1.1 and then explain how to modify it to the more general setting of Corollary 1.2.
@@
For $n\in\mathbb{N}$, consider the subgraph of $G$ induced by $\{\sigma=1\}\cap\Lambda_{n}(\rho)$ and denote its connected component containing $\rho$ by $\mathcal{C}_{n,\rho}$. Define $\mathsf{UST}_{n,G,\sigma}$ as the uniform spanning tree measure on $\mathcal{C}_{n,\rho}$. It is standard that the UST satisfies a negative correlation inequality, whence the limit of $\mathsf{UST}_{n,G,\sigma}$ as $n$ tends to infinity exists; see [9]. The limit is called free uniform spanning forest, we denote it by $\mathsf{FUSF}_{G,\sigma}$. Below we do not rely on [3, Theorem $5.13$] establishing that the FUSF is connected a.s. since that work assumes that $\rho$ has a finite expected degree.
@@
Take any finitary invariant bond percolation $\omega$ and let $T(G,\rho,\omega)$ be the number of trifurcation points of $\tau$ in $K_{\omega}(\rho)$. By the standard 3-partition argument of Burton and Keane (see [30, Lemma 8.5]), $T(G,\rho,\omega)$ is bounded above by $|\partial K_{\omega}(\rho)|$. Now set
@@ 4. Loop $\mathrm{O}(n)$ model: proof of Theorem 2
The graphical representation below is inspired by the Edwards–Sokal coupling [23] and the work of Chayes and Machta [15]. In a particular case of $n=2$, this representation was introduced by the second author and Spinka; it is described in [27].
@@
Let $n\in[1,2]$ and $x\in[1/\sqrt{2},1]$. Consider any translation-invariant Gibbs measure $\mathbb{P}$ for the loop $\mathrm{O}(n)$ model with edge-weight $x$. Sample a loop configuration $\omega$ from $\mathbb{P}$ and then $\xi$ from $\omega$. Then, $\xi$ exhibits no infinite connected component, almost surely.